2) Youve obtained the following balance sheet information for 2015 & 2016 along with income statement information for 2016 2016 $80 Catego 2015 $50 Cate 2016 $50 Sales Cost of Goods Sold Interest Expense Current Assets 420 60 $24 $5 Net Fixed Assets Accounts Payable Accrued Expenses Long-term Debt Owners Equit 250 30 25 150 95 40 250 150 Depreciation Tax Rate $6 30% a)Calculate the Operating Cash Flow for 2016 b)Calculate the Free Cash Flow for 2016 c) What is the free cash flow to creditors? What is the free cash flow to bondholders? d)If accounts payable increases, briefly explain in a sentence or two, how that would affect free cash flow for the company
Answer #1

OPERATING CASH FLOW FOR 2016:

A) Taxes = ( Sales - COGS - Depreciation - Interest)*Tax rate = (50-24-6-5)*30% = 4.5
Operating cash Flow = EBIT - Taxes + Depreciation = (Sales - COGS - Depreciation) -Taxes + Depreciation = (50-24-6) -4.5 +6 = 21.5

Free cash flow for 2016:

B) Free cash Flow = OCF - Increase in fixed Assets - Increase in Working capital = 21.5 - (420-250)-((80-100)-(50-55)) = = 21.5 -170 + 15 = -133.50
Free cashflow to creditors:


C) Free cash flow to Bond holders = Interest + Increase in debt = 5 +100 =105
Free Cash flow to creditors = Free Cash flow - Free Cash flow to bond holders = -133.50 -105 = -238.50

D) If accounts payable increases the free cash flow increases then cash will remain with the firm for longer time